# A Hybrid Machine Learning Framework for Predicting Hydrogen Storage Capacities: Unsupervised Feature Learning with Deep Neural Networks

### v4

@@ 5.3 Case II: Adopting the hybrid Approach
Performance evaluation was undertaken by computing the $R^{2}$ value across different latent space dimensions. Our observations revealed that the $R^{2}$ value exhibited an initial increase with the expansion of the latent space dimension, post which it fluctuated (Fig. 5) The peak $R^{2}$ value (0.85) was obtained with a latent space dimension of $z=8$ (Fig. 6).
@@ 6 Comparison with few other supervised models
In this section, we compare the performance of two supervised models, Lasso regression and Gradient Boosting regression, with the previously mentioned hybrid approach using an Autoencoder combined with a Multilayer Perceptron (MLP) 54, 55. We evaluate the models based on their R2 scores, which measure the proportion of the variance in the target variable that the models can explain. The results are shown in Fig.7. The R2 values obtained from the Lasso regression and Gradient Boosting regression are 0.48 and 0.8, respectively. These values indicate that both models have relatively lower predictive accuracy compared to the hybrid approach. The hybrid approach achieved an R2 score of 0.85, outperforming both linear models. These results suggest that the hybrid approach, combining an Autoencoder with an MLP, is more effective in capturing the underlying patterns and predicting the target variable compared to the individual linear models.
@@ 6.1 Identification of new hydrogen storage materials
To find new materials for efficient hydrogen storage, we used the following approach: We calculated the Pearson correlation function for different physically meaningful features, including the target - the storage capacity (Wt), which we show in Fig. 8 56. We take the weighted sum for all nine features and compute the correlation. It can be seen that the period number and electronegativity have the highest correlation with the weight percentage. A significant Pearson correlation between the period number and electronegativity with the target hydrogen storage capacity, suggests their pivotal roles in influencing a material’s ability to store hydrogen efficiently. The period number, reflecting the electron shell structure, directly impacts its hydrogen bonding capabilities, while electronegativity, indicative of an electron attraction strength of the material, influences the stability and capacity of hydrogen storage in metal hydrides. Therefore, we first scanned the metal hydrides from the Materials Project with these two features. Next, we constructed a 36-dimensional feature space using the above materials, as before (using four statistical measures for each one). Then, we used pre-trained embeddings from the autoencoder so that the property (hydrogen weight in percentage) can be computed using the MLP. Finally, we performed DFT-based calculations to re-verify these values.
@@
Figure 8: Correlation between different physically meaningful features and the target. Here atomic number ($Z$), period ($P$), electronegativity ($\chi$), electron affinity ($\chi_{a}$), atomic volume ($V_{a}$), atomic weight (W), fusion heat ($H_{f}$), and ionization energy ($IE$) is considered.
@@
Using the above approach we identified two materials Al11O18 and V2O5. Both the materials are stable from our formation energy calculations. The DFT calculated formation energies are respectively -3.34 eV/atom and -2.12 eV/atom for Al11O18 and V2O5. From the DFT approach the hydrogen storage capacities were calculated using the hydrogen adsorption energies 11 shown below,
$E_{ads}=\frac{1}{n}\times\left(E_{structure+nH}-\left(E_{structure}+\frac{n}{2}E_{H_{2}}\right)\right)$ (5)
@@
where Estructure+nH is the energy of the hydrogenated structure, Estructure and EH2 are the energies of the pristine structure and isolated H2, respectively, and n is the number of hydrogen atoms involved in the adsorption.
@@
To predict hydrogen storage capacity, all interstitial sites of Al11O18 and V2O5 are gradually filled with hydrogen, and the adoption energy is calculated 57. Figure 9 shows the change in adsorption energy with increasing hydrogen concentration. The adsorption energy increases as the hydrogen fraction increases. The maximum hydrogen fraction at which adsorption energy remains negative is used to predict hydrogen storage capacity. The predicted hydrogen storage capacities from the DFT calculations are respectively 4.61% and 3.83% while from the MLP model, we obtain values respectively 5.98% and 4.35%.
Figure 9: Adsorption energy of hydrogen in Al11O18 (a) and V2O5 (b) at different H concentration.
@@
In the Fig.10, we show the density of states (DOS) for the machine learning predicted materials and their hydrides for different hydrogen loadings. It was found that Al11O18 shows a more stable behavior when loaded with hydrogen compared to V2O5 because V2O5 relatively DOS at the Fermi energy for all the hydrogen loading. For both materials, H-s strongly hybridizes with the O-p states as can be seen from the projected DOS of Al11O18H4 and V2O5H shown in the down panel of the above figure. In the case of V2O5H, the anti-bonding states formed between V-d and H-s are also filled and therefore the system shows less stability.
@@ 7 Application to the Hydrogen Storage Materials Generated Using a Large Language Model
To show further the predictive capability of our approach, we apply it further to a set of materials that we generate using a large language model (LLM). Leveraging the power of GPT-2 (Generative Pre-trained Transformer version 2), a model renowned for its proficiency in natural language processing, we have generated a few new hydrogen storage materials. Unlike the resource-intensive traditional methods such as Generative Adversarial Networks (GANs) or diffusion models, which have limited testing in materials science, our approach utilizes a fine-tuned GPT-2 model, offering a more quick and targeted pathway to innovation. The details of the methodology are explained as detailed by Fu et. al 58 which leverages the advanced capabilities of LLMs originally developed for natural language processing, adapting them to the intricate language of material compositions. The fine-tuning process involves adapting a pre-trained GPT-2 model to the specific nuances of material science. This adaptation is facilitated by a custom vocabulary. which includes a comprehensive list of chemical elements and special tokens. This specialized vocabulary is crucial for accurately representing and processing material compositions. Our training process fine-tunes the model with a dataset specifically curated for hydrogen storage materials, allowing the model to learn and predict the complex patterns of material composition. The cornerstone of this methodology is the training of these models on comprehensive databases of known material compositions, including the ICSD (Inorganic Crystal Structure Database), OQMD (Open Quantum Materials Database), and Materials Project. This extensive assimilation of existing material data equips LLMs like GPT-2 with the ability to understand and replicate the intricate patterns and rules that govern material compositions. Through this process, the models are enabled to generate new, chemically plausible material compositions, which hold great potential for hydrogen storage applications. In an extension to the work by Fu et al. 58 on materials generation, we take a slightly modified approach where not only the chemical composition of materials but also a specific property: their hydrogen storage capacity is also considered to generate the chemical labels.
@@
As a result, the training of the model emphasizes the detailed relationship between a material’s composition and its storage capacity. Such integration enhances the model’s capacity to predict materials more accurately. The configuration of the model is tailored for this task, with an embedding size of 180, 2 attention heads, and 6 layers, which makes for a robust yet effective learning structure. Furthermore, the model’s vocabulary has been expanded to include numerical values, facilitating the inclusion of hydrogen storage capacity data.
Starting from a sequence of input tokens (comprising chemical symbols and storage capacity) the model first computes their embeddings. It then applies learned weight matrices to derive the usual queries, keys, and values matrices $Q$, $K$, and $V$, which are used to calculate the self-attention scores given by $\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ 59 where $d_{k}$ is the dimensionality of the key vectors. These scores determine how much each token in the sequence should attend to every other token, capturing both local and global dependencies within the sequence. During the training phase, the objective is to fine-tune a model’s parameters, denoted by $\theta$, to accurately predict material compositions from a given sequence of elements. This process involves presenting the model with sequences of tokens $[t_{1},t_{2},\ldots,t_{n}]$, each representing either a unique element within a material composition or a numerical value representing the storage capacity. The model learns by adjusting $\theta$ to maximize the log-likelihood of correctly predicting the next token $t_{n+1}$ in the sequence, formalized as $\max_{\theta}\sum\log P(t_{n+1}|t_{1},\ldots,t_{n};\theta)$. Here, $P(t_{n+1}|t_{1},\ldots,t_{n})$ represents the conditional probability of observing the next token $t_{n+1}$ given the preceding sequence $[t_{1},\ldots,t_{n}]$. The self-attention mechanism within the model facilitates this learning process by allowing each token to dynamically influence the prediction of the next token based on its context and relevance to the rest of the sequence.
@@
$)H2C2N$+6 -0.11 -3.5 -3.3 10.06 10.8 0.46 yes
[…]
@@
Here the critical hydrogen chemical potential, $\mu_{H}^{C}$ is defined as $\mu_{H}^{C}=1/n\left(E_{MAX+nH}-E_{MAX}\right)$, which is lower than the chemical potential $\mu_{H}$ of the hydrogen at gas phase at very low temperature (as our calculation corresponds to DFT temperature scale). ”FE” stands for the formation energy. From the above table, it is apparent that our model exhibits commendable performance and is quite reliable for the calculation of the hydrogen storage capacity. In the final table 2, we present an additional five materials that have been predicted by GPT-2 and whose hydrogen weight percentage has been computed by our model. As the atomic configurations of these materials remain elusive, we employ the methodology expounded by Kusaba et al 60, which employs the approach of metric learning, where the algorithm has been trained on a substantial amount of crystal structures that have already been identified to ascertain the isomorphism of crystal structures formed by two specified chemical compositions, yielding an accuracy of approximately 96.4%. For a given composition with an unknown crystal structure, this approach automatically selects a collection of template crystals from a database of crystal structures that possess nearly identical stable structures. When we subject the compositions predicted by GPT-2 to this method, we obtain a collection of structures for each composition. We consider the one with the highest spacegroup and then perform the structural optimization followed by the calculation of the storage capacity using our hybrid-ML approach. The rightmost column of the table visually represents the crystal structure of each material.
@@
B2NLi 7.18 166 [Uncaptioned image]
Table 2: Materials predicted by GPT-2 model and their storage capacities obtained using our hybrid ML approach. The spacegroup information was obtained via DFT optimization of the structure with the highest spacegroup predicted by the approach mentioned in Kusaba et al 60

### v5

@@ 5.3 Case II: Adopting the hybrid Approach
Performance evaluation was done by computing the $R^{2}$ value across different latent space dimensions. Our observations revealed that the $R^{2}$ value exhibited an initial increase with the expansion of the latent space dimension, post which it fluctuated (Fig. 5) The peak $R^{2}$ value (0.85) was obtained with a latent space dimension of $z=8$ (Fig. 6).
@@ 6 Comparison with few other supervised models
In this section, we compare the performance of two supervised models, Lasso regression and Gradient Boosting regression, with the previously mentioned hybrid approach using an Autoencoder combined with a Multilayer Perceptron (MLP) 59, 60. We evaluate the models based on their R2 scores, which measure the proportion of the variance in the target variable that the models can explain. The results are shown in Fig.7. The R2 values obtained from the Lasso regression and Gradient Boosting regression are 0.48 and 0.8, respectively. These values indicate that both models have relatively lower predictive accuracy compared to the hybrid approach. The hybrid approach achieved an R2 score of 0.85, outperforming both linear models. These results suggest that the hybrid approach, combining an Autoencoder with an MLP, is more effective in capturing the underlying patterns and predicting the target variable compared to the individual linear models.
@@ 6.1 Identification of new hydrogen storage materials
To find new materials for efficient hydrogen storage, we used the following approach: We calculated the Pearson correlation function for different physically meaningful features, including the target - the storage capacity (Wt), which we show in Fig.8 61. We take the weighted sum for all nine features and compute the correlation. It can be seen that the period number and electronegativity have the highest correlation with the weight percentage. A significant Pearson correlation between the period number and electronegativity with the target hydrogen storage capacity, suggests their pivotal roles in influencing a material’s ability to store hydrogen efficiently. The period number, reflecting the electron shell structure, directly impacts its hydrogen bonding capabilities. As the period number increases, so does the atomic radius. This can influence the hydrogen storage capacity because larger atoms can provide more interstitial sites for hydrogen atoms to occupy. On the other hand, electronegativity is a measure of the ability of an atom to attract and hold onto electrons. Elements with higher electronegativity tend to form more stable hydrides because they attract and bond more effectively with hydrogen atoms. Therefore, we first scanned the metal hydrides from the Materials Project with these two features. Next, we constructed a 36-dimensional feature space using the above materials, as before (using four statistical measures for each one). Then, we used pre-trained embeddings from the autoencoder so that the property (hydrogen weight in percentage) can be computed using the MLP. Finally, we performed DFT-based calculations to re-verify these values.
@@
Figure 8: Correlation between different physically meaningful features and the target (Wt). Here atomic weight (W), atomic number ($Z$), period ($P$), electronegativity ($\chi$), covalent radius ($r_{c}$), electron affinity ($\chi_{a}$), atomic volume ($V_{a}$), fusion heat ($H_{f}$), and ionization energy ($IE$) are considered.
@@
Using the above approach we identified two materials Al11O18 and V2O5. From our formation energy calculations, both materials are stable. The DFT calculated formation energies are respectively -3.34 eV/atom and -2.12 eV/atom for Al11O18 and V2O5. From the DFT approach the hydrogen storage capacities were calculated using the hydrogen absorption energies 12 shown below,
$E_{abs}=\frac{1}{n}\times\left(E_{structure+nH}-\left(E_{structure}+\frac{n}{2}E_{H_{2}}\right)\right)$ (5)
@@
where Estructure+nH is the energy of the hydrogenated structure, Estructure and EH2 are the energies of the pristine structure and isolated H2, respectively, and n is the number of hydrogen atoms involved in the absorption.
@@
To predict hydrogen storage capacity, all interstitial sites of Al11O18 and V2O5 are gradually filled with hydrogen, and the absorption energy is calculated 62. Figure 9 shows the change in absorption energy with increasing hydrogen concentration. The absorption energy increases as the hydrogen fraction increases. The maximum hydrogen fraction at which absorption energy remains negative is used to predict hydrogen storage capacity. The predicted hydrogen storage capacities from the DFT calculations are respectively 4.61% and 3.83% while from the MLP model, we obtain values respectively 5.98% and 4.35%.
Figure 9: Absorption energy of hydrogen in Al11O18 (a) and V2O5 (b) at different H concentration.
@@
In Fig.10, we show the density of states (DOS) for the machine learning predicted materials and their hydrides for different hydrogen loadings. It was found that Al11O18 shows a more stable behavior when loaded with hydrogen compared to V2O5 because V2O5 relatively DOS at the Fermi energy for all the hydrogen loading. For both materials, H-s strongly hybridizes with the O-p states as can be seen from the projected DOS of Al11O18H4 and V2O5H shown in the down panel of the above figure. In the case of V2O5H, the anti-bonding states formed between V-d and H-s are also filled and therefore the system shows less stability.
@@ 7 Application to the Hydrogen Storage Materials Generated Using a Large Language Model
To show further the predictive capability of our approach, we apply it further to a set of materials that we generate using a large language model (LLM). Leveraging the power of GPT-2 (Generative Pre-trained Transformer version 2), a model renowned for its proficiency in natural language processing, we have generated a few new hydrogen storage materials. Unlike the resource-intensive traditional methods such as Generative Adversarial Networks (GANs) or diffusion models, which have limited testing in materials science, our approach utilizes a fine-tuned GPT-2 model, offering a more quick and targeted pathway to innovation. The details of the methodology are explained as detailed by Fu et. al 44 which leverages the advanced capabilities of LLMs originally developed for natural language processing, adapting them to the intricate language of material compositions. The fine-tuning process involves adapting a pre-trained GPT-2 model to the specific nuances of material science. This adaptation is facilitated by a custom vocabulary. which includes a comprehensive list of chemical elements and special tokens. This specialized vocabulary is crucial for accurately representing and processing material compositions. Our training process fine-tunes the model with a dataset specifically curated for hydrogen storage materials, allowing the model to learn and predict the complex patterns of material composition. The cornerstone of this methodology is the training of these models on comprehensive databases of known material compositions, including the ICSD (Inorganic Crystal Structure Database), OQMD (Open Quantum Materials Database), and Materials Project. This substantial assimilation of existing material data enables LLMs such as GPT-2 to understand and recreate the complex patterns and laws that govern material compositions. Through this method, the models can produce new, chemically realistic material compositions with high promise for hydrogen storage applications. In an extension to the work by Fu et al. 44 on materials generation, we take a slightly modified approach where not only the chemical composition of materials but also a specific property: their hydrogen storage capacity is also considered to generate the chemical labels.
@@
As a result, the training of the model emphasizes the detailed relationship between a material’s composition and its storage capacity. Such integration enhances the model’s capacity to predict materials more accurately. The model’s configuration is tailored for this task, with an embedding size of 180, 2 attention heads, and 6 layers, making for a robust yet effective learning structure. Furthermore, the model’s vocabulary has been expanded to include numerical values, facilitating the inclusion of hydrogen storage capacity data. The GPT-2 transformer model training for generating new hydrogen storage materials was conducted using the Huggingface Transformers library 63.
The training and generation of materials using GPT-2 transformer model is shown in the Fig.12. Starting from a sequence of input tokens (comprising chemical symbols and hydrogen storage capacity) the model first computes their embeddings. The application of learned weight matrices is subsequently utilized to deduce the conventional queries, keys, and values matrices $Q$, $K$, and $V$, respectively. These matrices play a crucial role in the computation of self-attention scores, as expressed by $\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ 64, where $d_{k}$ represents the dimensionality of the key vectors. The evaluation of these scores dictates the level of attention that each token in the sequence should allocate to every other token, encompassing both local and global dependencies present in the sequence. During the training phase, the objective is to fine-tune a model’s parameters, say denoted by $\theta$, to accurately predict material compositions from a given sequence of elements. In this process, the model is presented with sequences of tokens $[t_{1},t_{2},\ldots,t_{n}]$, each representing either a unique element within a material composition or a numerical value representing the hydrogen storage capacity. The model learns by adjusting $\theta$ to maximize the log-likelihood of correctly predicting the next token $t_{n+1}$ in the sequence, formalized as $\max_{\theta}\sum\log P(t_{n+1}|t_{1},\ldots,t_{n};\theta)$. Here, $P(t_{n+1}|t_{1},\ldots,t_{n})$ represents the conditional probability of getting the next token $t_{n+1}$ given the preceding sequence $[t_{1},\ldots,t_{n}]$. By enabling every token to dynamically affect the prediction of the subsequent token based on its context and relevance to the remainder of the sequence, the self-attention mechanism in the model helps with this learning process.
@@
The model leverages a large corpus of known material compositions, learning intricate relationships and dependencies between different elements. By incorporating these learned patterns, the model can generate sequences that adhere to known chemical rules and potentially suggest novel compositions that exhibit desirable properties. This approach significantly accelerates the discovery process by providing a vast space of candidate materials for further computational or experimental validation.
To ensure that the generated compositions are meaningful, a few steps are further used. Initially, the sequences are cleaned to remove special tokens and ensure they conform to valid chemical notation. Next, they are filtered based on predefined criteria, such as limiting the number of unique elements and the total atom count, to focus on manageable and realistic compositions. In our generation process, we set the filtering criteria to ensure that the generated material compositions do not have more than 8 unique elements and the total number of atoms in any generated composition does not exceed 30. Furthermore, such generation and post-processing steps combined lead to the generation of sequences focusing on chemical compositions without storage capacities. Therefore, we generate only chemical compositions not the storage capacities though the training data includes such capacities.
@@
$\mathrm{TiAlN_{2}}$ 3 1.06 -4.8 -3.3 2.86 4.6 0.21 no
$\mathrm{V}_{2}\mathrm{H}_{2}$ 2+3 -0.2 -3.58 -3.3 2.83 3.24 0.29 no
$\mathrm{Zr}_{2}\mathrm{Ti}\mathrm{Al}$ 1 -0.34 -3.72 -3.3 0.4 1.1 0.13 yes
@@
Here the critical hydrogen chemical potential, $\mu_{H}^{C}$ is defined as $\mu_{H}^{C}=1/n\left(E_{MAX+nH}-E_{MAX}\right)$, which is lower than the chemical potential $\mu_{H}$ of the hydrogen at gas phase at very low temperature (as our calculation corresponds to DFT temperature scale). ”FE” stands for the formation energy. From the above table, it is apparent that our model exhibits commendable performance and is quite reliable for the calculation of the hydrogen storage capacity. In the final table 2, we present an additional five materials that have been predicted by GPT-2 and whose hydrogen weight percentage has been computed by our model. As the atomic configurations of these materials remain elusive, we employ the methodology expounded by Kusaba et al 65, which employs the approach of metric learning, where the algorithm has been trained on a substantial amount of crystal structures that have already been identified to ascertain the isomorphism of crystal structures formed by two specified chemical compositions, yielding an accuracy of approximately 96.4%. For a given composition with an unknown crystal structure, this approach automatically selects a collection of template crystals from a database of crystal structures that possess nearly identical stable structures. When we subject the compositions predicted by GPT-2 to this method, we obtain a collection of structures for each composition. We consider the one with the highest spacegroup and then perform the structural optimization followed by the calculation of the storage capacity using our hybrid-ML approach. The rightmost column of the table visually represents the crystal structure of each material.
@@
MgMnVTi 1.93 216 [Uncaptioned image]
Table 2: Materials predicted by GPT-2 model and their storage capacities obtained using our hybrid ML approach. The spacegroup information was obtained via DFT optimization of the structure with the highest spacegroup predicted by the approach mentioned in Kusaba et al 65